OpenStudy (anonymous):

An accident insurance company has reserves to cover claims on no more than 3 accidents. If all accidents have a probability of .01 of occurring, if all accidents are independent, and if the insurance company writes 100 policies, what is the probability that they will not have enough reserves?

OpenStudy (anonymous):

so if 4 accidents occur ,the insurance company wont be able to payout the accdnt insurance. You remember the expansion Robert shiller taught in econ 252 in his second class. go through it once again the probability = p^x * (1-p)^n-x * n!/ ( n-x)! p= probability of occurrence of an event 1-p=probability of an event not occuring n=total number of trials x=number of trials successful